OpenStudy (anonymous):

y=1/5x+3/5 y=4/7x-2 I got -7,2 but not sure how?

OpenStudy (zepp):

Since y = y, then \(\large \frac{1}{5}x+\frac{3}{5}=\frac{4}{7}x-2\), that would be the first step, right?

OpenStudy (anonymous):

yeah

OpenStudy (zepp):

Can you group similar term for me please? :D

OpenStudy (anonymous):

yeah

OpenStudy (zepp):

What do you get when you group them together?

OpenStudy (anonymous):

-4x-20y=-12, 4x+7y=-14

OpenStudy (anonymous):

-13y=-26 y=2 1/5x+(2)= 3/5 x=7

OpenStudy (zepp):

So you prefer the reduction method?

OpenStudy (anonymous):

yes

OpenStudy (zepp):

http://puu.sh/wfXe What you did was right, \(x=7\) :)

OpenStudy (anonymous):

okay thanks

OpenStudy (zepp):

Sorry for the delay, my internet is constantly cutting off :(

OpenStudy (anonymous):

that's okay